# Show-o: One Single Transformer to Unify Multimodal Understanding and Generation

### v6

@@ 5.1 Experimental Setup
Datasets. Three types of data are adopted for training Show-o: i) Text-only Data: We employ the publicly available RefinedWeb dataset (Penedo et al., 2023) to preserve the text reasoning capabilities of the pre-trained LLM. This dataset comprises approximately 1 billion instances (equivalent to 968 million individual web pages) and totals 2.8 terabytes of curated text data. ii) Image Data with Class Names: Show-o learns pixel dependencies using 1.28M images sourced from the ImageNet-1K (Deng et al., 2009) dataset. iii) Image-Text Data: For pre-training tasks correspond to multimodal understanding and generation, we assemble roughly 35M image-text pairs from the publicly available datasets including CC12M (Changpinyo et al., 2021), SA1B (Kirillov et al., 2023), and LAION-aesthetics-12M∗*∗∗*∗https://huggingface.co/datasets/dclure/laion-aesthetics-12m-umap. Additionally, we further increase the data scale to around 2.0B by incorporating DataComp (Gadre et al., 2024) and COYO700M (Byeon et al., 2022) with some filtering strategies. Note that, we employ ShareGPT4V (Chen et al., 2023) to re-caption these datasets. Additionally, LAION-aesthetics-12M and JourneyDB (Sun et al., 2023a) serve as high-quality datasets for the final fine-tuning. Following LLaVA-v1.5 (Liu et al., 2024b), we incorporate LLaVA-Pretrain-558K and LLaVA-v1.5-mix-665K for instruction tuning. Moreover, the GenHowTo dataset (Souček et al., 2024) is utilized for mixed-modality generation.
Evaluation Details. Following LLaVA (Liu et al., 2024b), we evaluate the multimodal understanding capabilities of Show-o on POPE, MME, Flickr30k, VQAv2, GQA, and MMMU benchmarks. Besides, following Stable Diffusion (Rombach et al., 2022), we adopt Fréchet Inception Distance (FID) on MSCOCO dataset to evaluate the potential generation fidelity of Show-o. Further, we follow SD3 (Esser et al., 2024) to evaluate the text-to-image generation capabilities of Show-o on the GenEval (Ghosh et al., 2023) benchmark. In experiments, we make comparisons with: i) understanding only models including LLaVA-v1.5 (Liu et al., 2024b), InstructBLIP (Dai et al., 2023), Qwen-VL-Chat (Bai et al., 2023), and mPLUG-Owl2 (Ye et al., 2024b); ii) generation only models including DALL·E (Ramesh et al., 2021), LDM (Rombach et al., 2022), DALL·E 2 (Ramesh et al., 2022a), SDv1.5 (Rombach et al., 2022), GigaGAN (Kang et al., 2023), PixArt (Chen et al., 2024), Imagen (Saharia et al., 2022), RAPHAEL Xue et al. (2024), LlamaGen (Sun et al., 2024), SDXL (Podell et al., 2023), and SD3 (Esser et al., 2024); unified models including Gemini (Anil et al., 2023), NExT-GPT (Wu et al., 2023b), Emu (Sun et al., 2023d), SEED-X (Ge et al., 2024), DreamLLM (Dong et al., 2024), CoDI (Tang et al., 2024), LWM (Liu et al., 2024a), and Chameleon (Team, 2024).
Implementation Details. We initially conduct joint training of Show-o using the RefinedWeb, a collection of 35M image-text pairs, and the ImageNet-1K for language modeling, image captioning, and class-conditional image generation, respectively, over 500K steps. Subsequently, we replace the class-conditional generation with the training for text-to-image generation using the 35M image-text pairs for an additional 1,000K steps. The base model is trained on 48 A100 (80GB) GPUs with a total batch size of 1,152. We employ the AdamW optimizer with a weight decay of 0.01, 5,000 steps of warm-up, and an initial learning rate of 1e-4 with a cosine scheduling. Finally, we fine-tune Show-o with filtered high-quality image-text pairs and adhere to the configuration of LLaVA-v1.5 for instruction data tuning. Note that, the current version of Show-o is based on Phi-1.5 (Li et al., 2023). In the following experiment sections, the default Show-o employs discrete image tokens as input for both multimodal understanding and generation. Show-o† and Show-o‡ indicate the use of continuous image representations from the pre-trained MAGVIT-v2 and CLIP-ViT (corresponding to options (b) and (c) in Fig. 5), respectively, for multimodal understanding and we discuss this exploration in Section 5.5.
@@ 5.2 Multimodal Understanding
5.2.1 Quantitative Evaluation
Table 1 presents the multimodal understanding capability of Show-o on public benchmarks, such as image captioning and visual question-answering tasks. i) The current version of Show-o is built upon Phi-1.5 (Li et al., 2023) and thus we follow LLaVA to train Show-o’s understanding only counterpart as our direct baseline, namely LLaVA-v1.5-Phi-1.5. The proposed Show-o exhibits comparable performance in all evaluation metrics to the baseline LLaVA-v1.5-Phi-1.5, which is dedicated and optimized to only multimodal understanding. This demonstrates the great potential of our framework to unify multimodal understanding and generation in one single transformer. ii) When comparing with understanding only models including InstructBLIP (Dai et al., 2023), Qwen-VL-Chat (Bai et al., 2023), and mPLUG-Owl2 (Ye et al., 2024b) on multimodal understanding, our model with a much smaller model size also achieves competitive performance on POPE, MME, Flickr30k and VQAv2 benchmarks and performs better on GQA benchmark. iii) Compared with unified models with a much larger number of parameters, such as NExT-GPT-13B (Wu et al., 2023b) and Chameleon-34B (Team, 2024), our model also achieves decent performance on Flickr30k benchmark and perform much better on VQAv2 benchmark. Given such promising results, we envision Show-o as a potential next-generation foundation model for unifying understanding and generation. These results also demonstrate the potential of scaling Show-o to achieve state-of-the-art performance.
5.2.2 Qualitative Results
We present Show-o’s visual question-answering capability and make comparisons with Chameleon and SEED-X in Fig. 7. It is evident that when presented with a query image, Show-o can describe the image in detail and respond to commonly asked questions, even addressing the unusual aspects within the image. In the example at the top of Fig. 7, Chameleon, SEED-X, and Show-o all provide comprehensive descriptions of the image’s main content. However, when asked, “Do you think the image is unusual or not”, Chameleon fails to correctly identify the unusual aspect, and SEED-X’s response, while identifying the unusual, lacks precision, stating “as living rooms are typically found on land”. In contrast, Show-o’s response, “as living rooms are usually indoors and designed for relaxation and entertainment”, is more accurate.
@@ 5.3 Visual Generation
5.3.1 Quantitative Evaluation
Results on MSCOCO 30K. We present zero-shot FID of Show-o on MSCOCO 30K in Table 2. It can be observed that, compared to generation models trained with larger numbers of parameters and training images such as GLIDE and DALL·E 2, Show-o achieves a better FID, i.e., 9.24, with only 1.3B parameters and 35M training data. Though GigaGAN, Imagen, and RAPHAEL obtain a relatively better performance than Show-o, they are much larger in model size (3B v.s. 1.3B) and trained with much more data. In comparison to unified models, Show-o also exhibits improvement. The above validates that Show-o, a unified transformer, can obtain competitive even better generation performance compared to individual models (generation only) with an equivalent or larger number of parameters and training data. However, it is worth noting that FID on MSCOCO 30K may not comprehensively accurate assessment of generation fidelity. The reason lies in the fact that existing generation models are commonly fine-tuned with high-quality and aesthetic images that do not align with the distribution of the MSCOCO dataset. Consequently, this mismatch leads to inaccurate measurement of generation fidelity.
@@ 5.3.1 Quantitative Evaluation
Show-o (Ours) | 1.3B | 0.98 | 0.80 | 0.66 | 0.84 | 0.31 | 0.50 | 0.68 |
Results on GenEval. Following SD3 (Esser et al., 2024), we evaluate text-to-image generation capability of Show-o on the GenEval benchmark (Ghosh et al., 2023) across six dimensions including “Single Object”, “Two Objects”, “Counting”, “Colors”, “Position”, “Color Attribute” and present the quantitative comparison in Table 3. One can observe that when comparing to the model in a similar size such as LDM (1.4B), Show-o obtains significantly better performance in all six metrics, with an improvement of around 0.31 overall. Besides, Show-o achieves a relatively better performance than DALL·E 2, which is 5 times larger in model size. Further, Show-o, with only 1.3B parameters, achieves even better performance than models with around a two-times larger number of parameters such as SD3 (2B). It indicates that our unified model’s generative capabilities are comparable to or even surpass those of specialized generation models. In comparison to unified models such as CoDI and SEED-X, Show-o also demonstrates significant improvements.
5.3.2 Qualitative Results
Qualitative Comparisons. We present qualitative comparisons with diffusion-based models, e.g., SDv1.5, SDXL, auto-regressive based model, i.e., LllamaGen, and unified models including LWM and SEED-X at the top of Fig. 8. One can observe that, given either short or long text prompts, Show-o can generate realistic images with consistent content described in the prompts. Compared to SDv1.5 and LlamaGen, Show-o demonstrates better visual quality and image-text alignment. For instance, as shown in the second column, both SDv1.5 and LlamaGen cannot fully comprehend the text prompt and miss some attributes such as sunset and blue domes in the generated images. In comparison to SDXL, Show-o exhibits comparable visual quality and alignment such as “a rally car race” and “stunning contrast against the vibrant sunset”. More samples generated by Show-o can be found at the bottom of Fig. 8. One can observe that Show-o is capable of generating diverse, interesting, and realistic visual content in a resolution of 512512. For example, Show-o can generate a futuristic style of car, a highly detailed face, cute objects, and vivid scenery with vibrant contrast.
Text-guided Inpainting and Extrapolation. As mentioned, Show-o naturally supports text-based inpainting and extrapolation without requiring any fine-tuning. We illustrate examples in Fig. 9. As shown on the top of the figure, given an input image and inpainting mask, Show-o can inpaint the original red trolley car to a blue sports car with sleek curves and tinted windows based on the user-provided text prompt. Besides, Show-o is capable of extrapolating the original image horizontally/vertically based on the given text prompt. Further, we can flexibly extrapolate the original image with new objects or scenes such as “red wildflowers” (as illustrated in the second row). One can observe that the pixels in both inpainted regions and extrapolated ones are consistent with the original ones. These cases significantly demonstrate the inherent advantages of Show-o over those auto-regressive models for downstream applications.
@@ 5.4 Mixed-Modality Generation of Video Keyframes and Captions
Here, we explore the mixed-modality generation ability of Show-o based on the text descriptions and video keyframes in the GenHowTo dataset. Given a sequence of interleaved text descriptions and video keyframes (as shown at the bottom of Fig. 4), Show-o is trained to predict the next text tokens or keyframe tokens conditioning on all preceding tokens. Thus, Show-o can generate mixed-modality of text descriptions and video keyframes. Examining a single frame, these tokens are generated in a diffusion manner. When considering the modeling of long sequences, as subsequent keyframes are produced based on all preceding text and image information, this can also be viewed as a form of temporal auto-regressive modeling. Consequently, it becomes feasible to generate consistent video keyframes continuously. Substituting video keyframes with video clips could offer a viable strategy for long video generation, a direction we aim to explore in the future.
We present qualitative examples in Fig. 10. As shown on the top of Fig. 10, given text prompts, Show-o can generate consistent video keyframes. Besides, we have tried to train Show-o using instructional examples. For example, given a question “Can you guide me through making Avocado and Apple Juice”, Show-o exhibits the capability to generate video keyframes with text descriptions related to the question. It is apparent that the generated keyframes are temporally consistent. This exploration reveals the potential of our model to be extended to the domain of long video generation, in which the model can produce a sequence of video clips. Specifically, the model continuously plans the next-scene textual prompt and proceeds to generate videos iteratively, allowing for the continuous generation of subsequent videos.
@@ 5.5 Ablation Studies
As illustrated in Fig. 5(a), the default Show-o adopts the pre-trained MAGVIT-v2 to tokenize input image to discrete tokens, which are then passed to the embedding layer to obtain embeddings as input for multimodal understanding. Beyond, we provide a systematic exploration of different design choices for the input of Show-o to enhance multimodal understanding. Specifically, as shown in Fig. 5(b) and (c), instead of discrete image tokens, we extract the continuous image representations from the pre-trained MAGVIT-v2 and CLIP-ViT, respectively, as input for Show-o when dealing with multimodal understanding. The experimental results are illustrated in Table 5. Through this exploration, we unveil the following lessons and insights.
Impact of Vision Encoder for Multimodal Understanding.

### v7

@@ 4.1 Experimental Setup
Datasets. We assemble two scales of datasets, i.e., around 35M and 2.0B image-text pairs, and collect around 2M high-quality data for multimodal understanding and generation fine-tuning. Besides, RefinedWeb (Penedo et al., 2023) is adopted as text corpora to maintain the language modeling capability. Appendix E provides more details about these datasets.
Evaluation Details. Following LLaVA (Liu et al., 2024b), we evaluate the multimodal understanding capabilities of Show-o on POPE, MME, Flickr30k, VQAv2, GQA, and MMMU benchmarks. Besides, we adopt Fréchet Inception Distance (FID) on MSCOCO dataset to evaluate the generation fidelity of Show-o. Further, we follow SD3 (Esser et al., 2024) to evaluate the text-to-image generation capabilities of Show-o on the GenEval (Ghosh et al., 2023) benchmark.
Implementation details. Current version of Show-o is based on Phi-1.5 (1.3B) (Li et al., 2023). In the following, the default Show-o employs discrete image tokens as input for both multimodal understanding and generation. Show-o† and Show-o‡ indicate the use of continuous image representations from the pre-trained MAGVIT-v2 and CLIP-ViT (corresponding to options (b) and (c) in Fig. 3), respectively, for multimodal understanding. Training details can be found in Appendix F.
@@ 4.2 Multimodal Understanding
Quantitative Evaluation. Table 1 presents the multimodal understanding capability of Show-o on public benchmarks, such as image captioning and visual question-answering tasks. i) The current version of Show-o is built upon Phi-1.5 and thus we follow LLaVA to train Show-o’s understanding only counterpart as our direct baseline, namely LLaVA-v1.5-Phi-1.5. The proposed Show-o exhibits comparable performance in all evaluation metrics to the baseline LLaVA-v1.5-Phi-1.5, which is dedicated and optimized to only multimodal understanding. This demonstrates the great potential of our framework to unify multimodal understanding and generation in one single transformer. ii) When comparing with understanding only models including InstructBLIP, Qwen-VL-Chat, and mPLUG-Owl2 on multimodal understanding, our model with a much smaller model size also achieves competitive performance on POPE, MME, Flickr30k and VQAv2 benchmarks and performs better on GQA benchmark. iii) Compared with unified models with a much larger number of parameters, such as NExT-GPT-13B and Chameleon-34B, our model also achieves decent performance on Flickr30k benchmark and performs much better on VQAv2 benchmark.
Qualitative Results. We present Show-o’s visual question-answering capability and make comparisons with Chameleon in Fig. 6. It is evident that when presented with a query image, Show-o can respond to commonly asked questions, even addressing the unusual aspects within the image. In the example of Fig. 6, when asked, “Do you think the image is unusual or not”, Chameleon fails to correctly identify the unusual aspect. In contrast, Show-o’s response, “as living rooms are usually indoors and designed for relaxation and entertainment”, is more accurate.
@@ 4.3 Visual Generation
Results on MSCOCO 30K. We present zero-shot FID of Show-o on MSCOCO 30K in Table 2. It can be observed that, compared to generation models trained with larger numbers of parameters and training images such as GLIDE and DALL·E 2, Show-o achieves a better FID, i.e., 9.24, with only 1.3B parameters and 35M training data. Though GigaGAN, Imagen, and RAPHAEL obtain a relatively better performance than Show-o, they are much larger in model size (3B v.s. 1.3B) and trained with much more data. In comparison to unified models, Show-o also exhibits improvement. However, it is worth noting that FID on MSCOCO 30K may not be a comprehensively accurate assessment of generation fidelity. The reason lies in the fact that existing generation models are commonly fine-tuned with high-quality and aesthetic images that do not align with the distribution of the MSCOCO dataset.
@@
| Show-o (Ours) | 1.3B | 0.98 | 0.80 | 0.66 | 0.84 | 0.31 | 0.50 | 0.68 | |
| Show-o‡ (Ours) | 1.3B | 0.98 | 0.85 | 0.67 | 0.81 | 0.28 | 0.55 | 0.69 |
Results on GenEval. One can observe in Table 3 that when comparing to the model in a similar size such as LDM (1.4B), Show-o obtains significantly better performance in all six metrics, with an improvement of around 0.24 overall. Besides, Show-o achieves a better performance than DALL·E 2, which is 5 times larger in model size, and SDXL. Further, Show-o, with only 1.3B parameters, achieves comparable performance to models with around two times larger number of parameters such as SD3 (2B). It indicates that our unified model’s generative capabilities are comparable to or even surpass those of specialized generation models. In comparison to unified models such as CoDI, SEED-X, and Chameleon, Show-o also demonstrates significant improvements.
Qualitative Results. We show image samples generated by Show-o in Fig. 7. One can observe that Show-o is capable of generating diverse, interesting, and realistic visual content in a resolution of 512512. For example, Show-o can generate a futuristic style of car, a highly detailed face, cute objects, and vivid scenery with vibrant contrast.
Text-guided Inpainting and Extrapolation. As mentioned, Show-o naturally supports text-based inpainting and extrapolation without requiring any fine-tuning. We illustrate examples in Fig. 8 (a). As shown on the top of the figure, given an input image and inpainting mask, Show-o can inpaint the original red trolley car to a blue sports car with sleek curves and tinted windows based on the user-provided text prompt. Specifically, we first tokenize the original image, mask those tokens to be inpainted, and then Show-o will gradually replace the masked tokens with predicted image tokens. Besides, Show-o is capable of extrapolating the original image horizontally/vertically based on the given text prompt. These cases significantly demonstrate the inherent advantages of Show-o over those autoregressive models for downstream applications.
@@ 4.4 Mixed-Modality Generation of Video Keyframes and Captions
Here, we explore the mixed-modality generation ability of Show-o based on the text descriptions and video keyframes in the GenHowTo dataset. Given a sequence of interleaved text descriptions and video keyframes (as shown at the bottom of Fig. 4), Show-o is trained to predict the next text tokens or keyframe tokens conditioning on all preceding tokens. Thus, Show-o can generate mixed-modality of text descriptions and video keyframes. Examining a single frame, these tokens are generated in a diffusion manner. When considering the modeling of long sequences, as subsequent keyframes are produced based on all preceding text and image information, this can also be viewed as a form of temporal auto-regressive modeling.
We have tried to train Show-o using instructional examples and present qualitative examples in Fig. 8 (b). For example, given a question “Can you guide me through making Avocado and Apple Juice”, Show-o exhibits the capability to generate video keyframes with text descriptions related to the question. It is apparent that the generated keyframes are temporally consistent.
4.5 Video Understanding and Generation
Beyond image understanding and generation, we have explored Show-o to support video understanding and generation. We inflate the existing pre-trained MAGVIT-v2 to support encoding videos into discrete tokens, compressing an 8 FPS video tensor of into . In this way, following the sequence format and omni-attention of multimodal understanding and generation introduced in Section 3.2, it is convenient to fine-tune the existing Show-o to involve video understanding and generation. One can observe visual examples in Figs. 8 (c) and (d) that Show-o can accurately comprehend and describe the variations in the video and generate consistent video frames with “a jeep car is approaching from a distance”. More examples can be found in Appendix H.
@@ 4.6 Ablation Studies
| # Image-text | Resolution | POPE | MME | Flickr30k | VQAv2 | GQA | MMMU |
| 35M | 2562 | 73.8 | 948.4 | 36.2 | 59.3 | 48.7 | 25.1 |
| 2.0B | 2562 | 76.2 | 1014.9 | 48.9 | 64.7 | 54.2 | 25.0 |
| 2.0B | 5122 | 80.0 | 1097.2 | 62.5 | 69.4 | 58.0 | 26.7 |
@@
As illustrated in Fig. 3(a), the default Show-o adopts the pre-trained MAGVIT-v2 to tokenize input image to discrete tokens, which are then passed to the embedding layer to obtain embeddings as input for multimodal understanding. Beyond, we provide a systematic exploration of different design choices for the input of Show-o to enhance multimodal understanding. Specifically, as shown in Fig. 3(b) and (c), instead of discrete image tokens, we extract the continuous image representations from the pre-trained MAGVIT-v2 and CLIP-ViT, respectively, as input for Show-o when dealing with multimodal understanding. We provide experimental results and insights in Appendix I.
Additionally, we present qualitative examples to illustrate the impact of sampling steps and classifier-free guidance for text-to-image generation in Appendix I. In observation, increasing the sampling steps can allow the synthesis of an image that closely adheres to the prompt and improve fidelity. Besides, the classifier-free guidance can significantly make the colors and contents more diverse and consistent with the given text prompt.
We also discuss the failure modes of Show-o in Appendix K.
5 Conclusion
6 Acknowledgments
This research is supported by the National Research Foundation, Singapore under its AI Singapore Programme (AISG Award No:AISG3-RP-2022-030).
We would like to express our sincere gratitude to Henry Hengyuan Zhao for his valuable discussions and insightful feedback on multimodal understanding, Mingrui Wang for his patient assistance in helping us set up the development environment, and Han Yao for providing preprocessed video datasets.
Appendix
Appendix A Preliminaries
where because of the Markov property.
the variational lower bound can be further expressed under the image distribution (referring to the proof provided by Zhang et al. (2024) as detailed in the Appendix B):
Appendix B Alternative Lower Bound for the variational diffusion
The constants and are:
The alternative lower bound can be derived as:
This proof is provided by Zhang et al. (2024).
Appendix C Training Pipeline
i) Image Token Embedding and Pixel Dependency Learning: We employ RefinedWeb (Penedo et al., 2023) dataset to train Show-o to maintain the language modeling ability. Meanwhile, ImageNet-1K dataset (Deng et al., 2009) and large-scale image-text pairs are adopted to train Show-o for class-conditional image generation and image captioning, respectively. Here, we directly leverage the class names from ImageNet-1K as textual inputs for learning class-conditional image generation. This stage primarily involves the learning of new learnable embeddings for discrete image tokens, pixel dependency for image generation, and alignment between image and text for image captioning.
Appendix D Inference details
In inference, two types of predictions, i.e., text and image tokens, are involved in Show-o. In multimodal understanding, given the conditional image and questions, text tokens are auto-regressively sampled from the predicted tokens with higher confidence. In visual generation, given text tokens and tokens as initial input, Show-o predict logits in parallel, where and is the time step. Following the work (Chang et al., 2023), we compute both the conditional logit and unconditional logit for masked tokens. The final logit of each token is obtained by the following equation with a guidance scale :
For each token at location , we sample an image token from the codebook based on the predicted probability and indicate its predicted score as the confidence of this sampled token. The confidence of the unmasked token is set as 1.0. Next, we compute the number of image tokens that should be re-masked based on the mask scheduling function , where . More details of mask scheduling functions can be found in MaskGIT (Chang et al., 2022). Subsequently, we replace the predicted image tokens with token based on the following metric:
Appendix E Dataset details
[…]
Appendix F Implementation Details
We initially conduct joint training of Show-o using the RefinedWeb, a collection of image-text pairs, and the ImageNet-1K for language modeling, image captioning, and class-conditional image generation, respectively, over 500K steps. Subsequently, we replace the class-conditional generation with the training for text-to-image generation using the around 35M image-text pairs for an additional 1,000K steps. The base model is trained on 48 A100 (80GB) GPUs with a total batch size of 1,152. We employ the AdamW optimizer with a weight decay of 0.01, 5,000 steps of warm-up, and an initial learning rate of 1e-4 with a cosine scheduling. Finally, we fine-tune Show-o with around 1M internal high-quality image-text pairs and adhere to the configuration of LLaVA-v1.5 for instruction data tuning. Note that, the current version of Show-o is based on Phi-1.5 (Li et al., 2023). In the following experiment sections, the default Show-o employs discrete image tokens as input for both multimodal understanding and generation. Show-o† and Show-o‡ indicate the use of continuous image representations from the pre-trained MAGVIT-v2 and CLIP-ViT (corresponding to options (b) and (c) in Fig. 3), respectively, for multimodal understanding and we discuss this exploration in Section 4.6.
Appendix G Text Prompts
“A 3D render of a futuristic car made of glass, driving through a city of mirrors.”
“A colorful cartoon of a tiger camouflaged in an abstract art painting, its stripes merging with the wild brushstrokes. ”
“The image features a stylized stained glass illustration of a hummingbird with vibrant colors, set against a backdrop of swirling patterns and a large sun. The composition includes floral elements and intricate details, creating a vivid and dynamic scene that emphasizes the beauty of the bird. The colors range from greens to reds, enhancing the lively and artistic aesthetic of the piece. ”
“A 3D render of a surreal explosion scene on the shore of a beautiful white sand beach with crystal clear water. The explosion has a spatter of oil paint with pastel colors and a thick consistency. The explosion is in a quiet and serene environment. A beautiful Japanese woman with a dress compacted to the sea is seen. There are butterfly petals and flowers with an ethereal glow and bioluminescence. There are pink and blue roses, and the overall image has a surreal and dreamlike quality. ”
“A hyper-realistic close-up photograph of a woman’s face, focusing on the left side. The image is highly detailed and realistic, showing voluminous glossy lips slightly parted, a well-defined nose, and open eyes with long eyelashes that cast shadows on the skin. The eye color is crystal clear almond green. The skin texture is crisp, with incredible detail of natural, lush skin and pores and freckles, with subtle highlights and shadows that give a realistic, close-up appearance. ”
“A 3D render of a cute, round rice ball character named Mochi, with big, sparkling eyes that convey curiosity and joy. Its body is a soft, fluffy white with a slight sheen, resembling freshly cooked rice. Mochi has small, rosy cheeks that give it a warm, friendly expression. A tiny smile brightens its face, and it often sports a colorful ribbon tied around its ”waist,” adding a playful touch. Mochi’s arms and feet are cartoonishly short, allowing it to bounce adorably around its surroundings. This time, Mochi is placed against a background that is a vibrant explosion of colors, with bright hues of fuchsia, turquoise, lemon yellow, and emerald green creating a canvas of vibrant contrasts and playful energy. The clashing colors make Mochi’s soft white body and rosy cheeks stand out even more, inviting viewers into a world of cheerful exuberance and visual delight.”
Appendix H More Examples of Video
We provide more examples of video understanding and generation in Fig. 10.
Appendix I Ablation Studies